Electronic Notes in Theoretical Computer Science 5 (1996) to appear

An Example Paper

My Conference Organizers^{1}

My Department My University My City, My Country

My Co-author

My Co-author’s Department My Co-author’s University My Co-author’s City, My Co-author’s Country

E. N. T. C. S.

Elsevier Science B. V.

Abstract This is a short example to show the basics of using the entcs style macro files.

1 Introduction

This short note provides a guide to using the entcs macro package for preparing papers for publication in your conference Proceedings. The Proceedings may be printed and hard copies distributed to participants at the meeting; this is an option to Confereence Organizers may choose to exercise. The Proceedings also will be a volume in the new series Electronic Notes in Theoretical Computer Science (ENCS), which is published under the auspices of Elsevier Science B. V., the publishers of Theoretical Computer Science.

The entcs macro package consists of three files:

entcs.sty, the basic style file, entcs12.sty, a file containing information for typesetting papers at 12pt,

and entcsmacro.sty, a macro file containing the definitions of some of the theorem-

like environments and a few other tidbits.

1 Thanks to everyone who should be thanked

\begin{matrix}c \\ ◯ \end{matrix} 1996 Elsevier Science B. V.

My Lastname

The file entcs12.sty is called automatically by entcs.sty. Additional macro files can be added as arguments to \documentstyle. The file entcsmacro.sty must be included in the list, as is done at the start of the source file for this paper.

The formatting these style files impose should not be altered – the reason for using them is to attain a uniform format for all papers in the Proceedings of which your paper is a part.

The file instraut.dvi contains information about the macro package

elsart.sty and elsart12.sty, from which the files entcs.sty and entcs12.sty

are derived. Reading this file should answer most of the basic questions that might arise. It also gives some basic rules about using the macro package.

2 Frontmatter

The biggest difference between a “usual” L^{A}TEX style such as article.sty and the entcs package is that the entcs macro package requires the title, author’s name or names, abstract and “thanks” all to be included within the frontmatter environment. At the beginning of the source file for this paper, you’ll notice this. Also, you’ll notice that the usual \maketitle is absent; it no longer is needed. The entcs style package automatically generates the title, author’s name and address, and related material at the beginning of the paper.

The entcs macro package provides two alternatives to listing authors names and addresses. These are described in detail in the file instraut.dvi. Basically, listing each author and his or her address in turn, is the simplest method. But, if there are several authors and two or more share the same address (but not all authors are at this address), then the method of listing authors first, and then the addresses, and of referencing addresses to authors should be used.

Also, notice that acknowledgment of support (the contents of \thanks) should be done by a separate listing of \thanks[NSF]{To the NSF} with the optional argument – [NSF] – being used for \thanksref which is attached to those authors acknowledging such support. It is important that the \thanks not be included within the scope of \author{} or of \title{}, but it must be within the scope of the environment frontmatter.

More details about added terms such as \collab can be found in inst,dvi, if they are needed.

Also, notice that the command \lastname{My Lastname}} has been included before the frontmatter begins. This command should contain the last names of the authors of the paper. If there are no more than three authors, then they should be listed with the word “and” between the last two; if more than three authors collaborated on the paper, then the first author only should be listed, together with {\em et al}. This command creates the headline for each page after page 1.

Finally, please be sure to include an abstract for your paper.

3 Sectioning and Environments

Since ENTCS will be published through the auspices of Elsevier Science B. V., their style files have been used to create end entcs macro package. To illustrate that this package is not much different than most of the ones one encounters, we have the following.

Definition 3.1 A file is derived from another if it is obtained with only a few modifications from the original file.

Theorem 3.2 Both of the files entcs.sty and entcs12.sty are derived from

elsart.sty and elsasrt12.sty, respectively.

Proof. This is clear from the definition. □

If one wants to start a proof with a descriptive word, such as “sketch”, then one can use the \begin{proof*}...\end{proof*} environment, as in

Proof (Sketch) This is clear from the definition. □

The main differences between files entcs.sty and entcs12.sty and the ones created by Elsevier are naming of the series Electronic Notes in Theoretical Computer Science, which is where papers prepared using these style files normally would appear.

In any case, all of the usual features of L^{A}TEX are available with these style files – it is only the formatting that has been rigorously defined. Thus, one has available the sectioning commands \section, \subsection, \paragraph and \subparagraph. The numbering scheme used is one under which Theorem 1.2.3 is the third numbered item in second subsection of the first section of the paper. In order to facilitate cross-references, all of the named environments given below are numbered, and all use the same number scheme.

The file entcsmacro.sty contains additional information that is needed to typesest a paper. It also has the definitions of the AMS euler and blackboard bold fonts included. If you have these fonts available on your system, then you can access them by uncommenting the lines indicated in the file entcsmacro.sty. If you want to use symbols for the natural numbers, the reals, etc., then we prefer that you use the blackboard bold fonts, and not plain bold fonts.

Also, the names of theorem-like environments are provided in entcsmacro.sty. With the exception of the environment Algorithm, the names of all of these are full name, rather than a shortened version. The environments provided and their names are

• \begin{theorem} ... \end{theorem} for Theorems,
• \begin{lemma} ... \end{lemma} for Lemmas,
• \begin{corollary} ... \end{corollary} for Corollaries,
• \begin{proposition} ... \end{proposition} for Propositions,
• \begin{criterion} ... \end{criterion} for Criteria,
• \begin{alg} ... \end{alg} for Algorithms,

• \begin{definition} ... \end{definition} for Definitions,
• \begin{conjecture} ... \end{conjecture} for Conjectures,
• \begin{example} ... \end{example} for Examples,
• \begin{problem} ... \end{problem} for Problems,
• \begin{remark} ... \end{remark} for Remarks,
• \begin{note} ... \end{note} for Notes,
• \begin{claim} ... \end{claim} for Claims,
• \begin{summary} ... \end{summary} for Summary,
• \begin{case} ... \end{case} for Cases, and
• \begin{ack} ... \end{ack} for Acknowledgements.

For example,

Algorithm 1 Step 1: Write the paper Step 2: Format it with the entcs macro package Step 3: Ship the whole thing to Mike

4 References

All the cross-referencing facilities of L^{A}TEX are supported, so one can use \ref{} and \cite{} for cross-references within the paper and for references to bibliographic items. As is done in this note, the References section can be

composed with \begin{thebibliography}...\end{thebibliography}. Al-

ternatively, BibTEX can be used to compile the bibliography. Whichever one is used, the references are to be numbered consecutively, rather than by authordefined acronyms. Of course you can use your own acronyms for easy reference to each of the items in the bibliography, as has been done with the listing for this short note.

However, note that the references should not be started with a new \section command.

5 Summary

The entcs macro package should be easy to use and should provide a uniform layout for all the papers in the Proceedings.

Problem 5.1 Finish the paper and get it to your Program Chairman on time!

When you have finished preparing your paper, send a copy of the source file, together with any macro files that are needed to your Program Chairman. If the files are extensive, you can place copies in the pub/incoming sub-directory of the ftp directory on the machine indicated by your Program Chairman using anonymous ftp. If you do this, please send me email to alert me that the file(s) are here.

References

[1] Civin, P., and B. Yood, Involutions on Banach algebras, Pacific J. Math. 9 (1959), 415–436

[2] Clifford, A. H., and G. B. Preston, “The algebraic theory of semigroups,” Math. Surveys 7, Amer. Math. Soc., Providence, R.I., 1961 (vol. 1)

[3] Drazin, M. P., Regular semigroups with an involution, Proc. of a Symposium on regular semigroups, N. Illinois University, 1959

[4] Easdown, D., and W. D. Munn, On semigroups with involution, Bull. Austral. Math. Soc. 48 (1993), 93–100

[5] Easdown, D., and W. D. Munn, Trace functions on inverse semigroup algebras, U. of Glasgow, Dept. of Math., preprint 93/52

[6] Munn, W. D., Semiprimitivity of inverse semigroup algebras, Proc. Roy. Soc. Edinburgh Ser. A 93 (1982), 83–98

[7] Munn, W. D., Some remarks on real and complex algebras, U. of Glasgow, Dept. of Math., preprint 93/70

[8] Passman, D. S., “The algebraic theory of group rings,” Wiley-Interscience, New York, 1977

[9] Shehadah, A. A., “Embedding theorems for semigroups with involution, “ Ph.D. thesis, Purdue University, Indiana, 1982

[10] Weyl, H., “The classical groups,” 2nd Ed., Princeton U. Press, Princeton, N.J., 1946